Adjective or Noun?

Lajos Marosán

§1 In ¶106 of A New English Grammar, Henry Sweet (1900) says that in the phrases silk thread, gold watch the words silk, gold, though they are similar to adjectives in that they are used as attribute words very much as the adjective silken, they are not adjectives, but remain nouns — i.e. we cannot talk about conversion in such cases (in ¶105ff. Sweet discusses conversion of the parts of speech) — because the words in question do not admit of any of the grammatical features characteristic of adjectives, such as premodification by very, comparison: *very silk, *more silk, whereas it is possible to say very silken, more silken, which proves that silken is an adjective. He adds that more silk is a grammatical phrase and suggests the idea of ‘a larger quantity of silk’, which meaning is totally different from the one under discussion. To put it in a more current terminology, Sweet suggests that the function of premodifier in a NP can be realised by different lexical categories, viz. noun or adjective.

50 Lajos Marosán

A similar approach is advocated by Quirk et al. (1985). They say if nouns are used attributively, they are superficially similar to peripheral adjectives in satisfying criterion (a), i.e. the potential to occur in prenominal or attributive position (op.cit. : 410). These premodifying nouns, however, do not satisfy the other criteria for adjectivehood, viz. (b) predicative use after the copula seem, (c) premodification by very, and (d) comparison.^{1} In English, then, the following structures are ungrammatical: *The station seems bus; *a very bus station; *a busser station, even if a bus station is absolutely correct. Again, it seems Quirk et al. distinguish between form and function, and seem to suggest that the fact that both nouns and adjectives may occur in the same function does not mean that they also belong to the same lexical category. Elsewhere Quirk et al. discuss the problem of conversion (1985 : 1562). A word is considered as “a noun converted into an adjective” if it also occurs in predicative position, i.e. can appear after the copula be (which is not exactly the same criterion as (b),^{1} where occurrence after seem was crucial for distinguishing between adjectives and other word classes). Only adjectives can occur after the copula seem; nominal predicates are usually preceded by the verb be, still, they consider the following sentences grammatical: a brick garage ∼ the garage is brick; reproduction furniture ∼ this furniture is reproduction; Worcester porcelain ∼ this porcelain is Worcester. Though brick, reproduction and Worcester are full-fledged adjectives, they are not gradable, and only marginally admit of premodification.

Curme (1935 : 42ff.) seems to give a different answer to the same question, i.e. whether a prenominal noun is a noun or an adjective. He analyses boy in a boy actor as a noun used as an adjective. Huddleston writes that

“… this type of locution, ‘an X used as a Y’, with X and Y standing for different parts of speech, is extremely common in traditional grammar; it is, however, quite unclear and in need of explication. Is it being claimed that an X used as a Y actually is a Y or that it is merely functionally like a Y? If it actually is a Y, then in what sense is it also an X? … ” (1984 : 93)

As far as I understand Curme, one of the questions, at least, can be answered: Curme himself offers the explication Huddleston longs for. He says that while both nouns and adjectives are used in premodification, there are pairs with slightly different meaning. E.g. a boy actor has a concrete, whereas a boyish actor an abstract meaning, because “The new adjective is concrete, the older form abstract” (1935 : 44). This is a fairly straightforward statement: Curme seems to say that boy in a boy actor is an adjective. However, I find it problematic to apply the same principle to “adverbs, prepositional phrases, sentences used as adjectives” (op.cit. : 43). I also wonder whether ‘X has the force of Y’ is synonymous with ‘X is used as Y’. In this article I want to believe that Curme’s position i.e. a prenominal noun is actually an adjective is basically correct, though one which requires some clarification, not necessarily in Curme’s spirit.^{2}

§2 It is interesting that authors like Huddleston and Quirk et al. should find it difficult to accept that a noun, in certain positions, can be converted into an adjective, or becomes an adjective; or to put it in a slightly different, but a more non-committing way, that the same word can be a noun in certain contexts and adjective in others.^{3} This is all the more interesting since the same authors state or at least seem to say, that a word such as all is a predeterminer in the phrase all the girls, and it is a pronoun in All were present. Or that is a central determiner in that book, a pronoun in Look at that and an adverb in that beautiful (see Huddleston 1984, Chapters 6 & 7; Quirk et al. 1985, Chapters 5 & 6). Adjective–adverb homomorphs are also similar (Quirk et al. 1985 : 405) fast car/he drives fast. The same authors discuss (op.cit. : 288) reclassification of proper nouns as common nouns in special circumstances. I wonder whether prenominal position is special enough for nouns to be reclassified as adjectives. Whether or not such cases are subsumed under ‘conversion’ by these authors (cf. Sweet ¶105ff.) is immaterial. It would be interesting to know why Huddleston and Quirk et al. accept conversion — let me apply Sweet’s term to these cases—in certain contexts, but not in others.

§3 Quirk et al. (1985 : 410) contrast the NPs the large station and the bus station to argue for their position, i.e. that large is an adjective functioning as premodifier, and bus, which also functions as premodifier, is a noun. In other words, both words keep their identities, i.e. class membership, even if they occur in the same function. They argue that large is capable to occur in different positions, and inflects for comparison, so it satisfies all the four criteria (op.cit. : 404) of adjectivehood; i.e. (a) a large station,

(b) The station seems large, (c) a very large station and (d) a larger station. They go on to show that, since bus does not admit of any of the above features, it definitely cannot be an adjective. Their arguments are fair and acceptable but I am afraid they are not on the right track. I do not think that a Noun premodifier+N structure, and an Adjective premodifier+N are parallel the way it is presented by Quirk et al. An adjective like large is not the right one. On the basis of its semantic relation to the head noun, it is called a qualitative adjective, and as such, it can be easily shown that it is an adjective, moreover, a central member of its class (see op.cit. : 404 again). In my opinion, what are traditionally called ‘classifying adjectives’ are the real partners of nouns in premodification.^{4} A typical classifying adjective is just as insensitive to the criteria for adjectivehood as nouns, i.e. they are not gradable, cannot be premodified (in the same way, at least) and they usually do not occur in predicative position. How do we know, then, that a word like nuclear is really an adjective? The only criterion it satisfies is (a), i.e. it occurs attributively (just like nouns). Classifying adjectives are similar to prenominal nouns in that they identify someone or something as a member of a class. They are not gradable, because the thing or person is either the member of the class or not: financial help, abdominal pain (Sinclair 1990 : 66), nuclear scientist (Ferris 1993 : 22). Since adjectives which are confined to attributive position are insensitive to the inflectional and syntactic tests offered by Quirk et al., therefore, there might be other reasons why we want to see these words classed as adjectives. Some authors even consider “some” of the attributive adjectives as postdeterminers (in Quirk et al.’s sense; see Chapter 5). Halliday states that “The post-Deictic identifies a subset of the class of ‘thing’ … ” (1985 : 162). Words such as other, customary, habitual, normal, etc. belong here. On the other hand, a noun in premodification is insensitive to tests which could prove that it is a noun. In a bus station, bus cannot be pluralised, does not take determiners, and does not inflect for the genitive case. If we destroy the N+N structure and move bus into other positions, it automatically starts behaving as a normal noun: These buses are pink; The pink buses are always late; This is a bus; some buses; a lot of buses, etc. (The genitive is, as far as I can judge it, incompatible with bus anyway.) In a trouser pocket; scissor movement (Huddleston 1984 : 135), it looks as if these plural nouns wanted to shed whatever is nouny (Wierzbicka’s expression; 1986) in them. Unfortunately, there are examples where the plural marker is retained, such as parks commissioner, wages clerk (Sadler & Arnold 1994 : 218); sports car, etc.

§4 Coordination facts might offer evidence which provides syntactic support for my claim about prenominal nouns. It seems that only identical categories can be conjoined: I met your father and mother (Radford 1988 : 75); John is very kind and considerate (op.cit. : 76). This, however, does not explain why the following structures sound odd: A gamekeeper and a gun wounded him; a new and financial company (Ferris 1993 : 127). In the first instance there are two noun phrases conjoined, in the second two adjective phrases,^{5} so in both cases we have identical categories conjoined, still their coordination is strange, because the conjoint constituents only share the same phrasal and lexical, but not semantic properties. In the first example, the noun phrase a gamekeeper plays the role of agent, whereas a gun is instrument. In the second example, though we have two representatives of the same phrasal category, AP, the fact that new is a qualitative and financial is a classifying — or in Ferris’s terminology (1993) associative — adjective makes the structure odd, if not hopelessly ungrammatical. It is more usual to conjoin adjectives belonging to the one or the other category: marital and legal advisors; young and old citizens (op.cit. : 135).^{6} Therefore, it is not surprising that it is possible to conjoin an adjective and a noun when both are classifiers, like in They detest both suburban and city life; She likes both woollen and cotton dresses (Quirk et al. 1562 fn.). I would like to believe that the data above at least motivate the claim that city and cotton are adjectives in the NPs they occur in. I am aware that coordination evidence is far from conclusive: it is possible to use it to show that two different categories may share the same function, but it is not impossible to state on the basis of the above data that two elements which have been believed to belong to two separate categories can, in fact, be subsumed under the same.^{7}

§5 Jespersen in The Philosophy of Grammar says that substantives (i.e. nouns) are more special than adjectives, they are applicable to fewer objects (1992 : 74ff.). Nouns are also more complex, i.e. they suggest many distinguishing qualities, whereas an adjective singles out one quality; “in the parlance of logicians, the extension of a substantive is less, and its intension is greater than that of an adjective” (op.cit. : 75). He illustrates his point by invoking words “which may be used in both capacities”. The meaning of substantivised adjectives such as cathedral, the blacks, evergreens (op.cit. : 76) are more special than that of their corresponding adjectives. He also finds evidence in nouns made into adjectives, whose meanings are less special than that of the substantive of which they were formed. E.g. in French rose, mauve, puce (op.cit. : 76) are more general as colour adjectives than as nouns. Some of the examples Jespersen illustrates his point with have the structure Adjective+Noun. I would like to apply his statement to Premodifier+Noun structures in general, i.e. in my opinion, Jespersen’s arguments, if valid at all, are true of nominal noun phrase premodifiers, and not adjectives only. So I want to say that a premodifier, be it an adjective or a noun, denotes a quality, therefore, is less special than its head noun. So if a noun functions as premodifier, it loses its nouniness, its specialised character, and the properties and characteristics a noun has. By entering into grammatical relationship with another noun, it changes its ‘cognitive structure’ in Hewson’s sense (1991). Boy, then, in the noun phrase a boy actor only denotes one quality, or characteristic, similar to the nouns which are made into adjectives (see Jespersen’s examples above). From a semantic point of view, then, a prenominal noun is ‘poorer’ than when it functions as head; but this can only be seen if the construction is kept intact. If we examine a prenominal noun separately, it offers itself as a noun proper to our scrutiny. At the moment, however, I do not think I can offer a description or explanation of what type of quality or characteristic the ‘adjective’ boy has.^{8}

§6 In this article I have reconsidered the status of prenominal nouns, and tried to argue that my intuition that such nouns are actually adjectives is correct. Whether these words are adjectives or nouns is not so important, in my opinion, as the discussion — which leads us to the conclusion that this or that word is an adjective or a noun — of the relationship between a premodifier and its head — whatever lexical category we claim them to be—, and the arguments that can be brought up for or against one or the other solution, and at last, what type of evidence is relevant and conclusive. There are some points to be cleared up: (a) what meaning these nounadjectives have, i.e. it should be explained what property or characteristic the ‘noun’ is left with when it is reclassified as an adjective in attributive position; (b) whether all other form of premodification should be classed as adjectives; (c) whether these adjectives can precede, and on what conditions, the pro-N-bar one, like in an American, not the john one, poised his camera to shoot.^{9}

notes

[1] See page 404.

[2] In general I agree with Huddleston’s criticism.

[3] See Halliday & Hasan (1976 : 37ff.) An item is put into a semantic category e.g. this is classified as ‘selective’, and can have two grammatical functions: modifier and head, in different contexts.

[4] Halliday talks about Classifiers, subsuming under this label nouns, adjectives and verbs (=participles), which occur prenominally and indicate “a particular subclass of the thing in question; e.g. electric trains, passenger trains, a stopping train, a tied note” (1984 : 164). Sadler & Arnold (1994) call these property modifiers. See Zwicky (1985) for a different interpretation of “a particular subclass of the thing”. He says that those penguins are a type of penguins, will leave a kind of leaving; according to his views, then, a new company would be a type of company, but so would a financial company. I agree with those authors who say that new and financial are related to their respective head nouns in a different way.

[5] See Sadler & Arnold (1994) for a different analysis of the APs.

[6] Sadler & Arnold write that nominal (=classifying) adjectives do not coordinate: *solar and lunar heat (1994 : 210). It seems, however, that they do not object to coordinated prenominal nouns: kitchen and bathroom furniture (op.cit. : 198 fn.3)

[7] Curme might have had the same reasons to say that whatever occurs in a position an adjective can have is actually an adjective — it is not at all clear to me either whether he identifies the parts of speech with the typical function(s) they perform.

[8] See Wierzbicka’s criticism (1986) of Jespersen. She also suggests that “phrases such as child-bride or gorod-geroj (‘city hero’, Russian) evoke two superimposed

Adjective or Noun? 55

images rather than one complex image, because each noun preserves its own basic categorisation” (op.cit. : 371).

[9] Quirk et al. (1985 : 1244) quote from Anthony Burgess’s Inside Mr Enderby.

references

Curme, G. O. (1935) Parts of Speech and Accidence. Boston: Heath. Ferris, C. (1993) The Meaning of Syntax. A Study in the Adjectives of English. London:

Longman. Halliday, M. A. K. (1985) An Introduction to Functional Grammar. London: Edward

Arnold. Halliday, M. A. K. and R. Hasan (1976) Cohesion in English. London: Longman. Hewson, J. (1991) ‘Determiners as heads’. Cognitive Linguistics 2. 317–337. Huddleston, R. (1984) Introduction to the Grammar of English. Cambridge: CUP. Jespersen, O. (1992) The Philosophy of Grammar. Chicago: The University of Chicago

Press. Quirk, R., S. Greenbaum, G. Leech and J. Svartvik (1985) A Comprehensive Grammar

of English. London: Longman. Radford, A. (1988) Transformational Grammar. Cambridge: CUP. Sadler, L. and J. D. Arnold (1994) ‘Prenominal adjectives and the phrasal/lexical distinc-

tion.’ Journal of Linguistics 21. 187–227. Sinclair, J., ed. (1990) Collins Cobuild English Grammar. London: Harper Collins. Sweet, H. (1900) A New English Grammar. Oxford: Clarendon. Wierzbicka, A. (1986) ‘What is in a noun?’. Studies in Language 12. 353–375. Zwicky, A. (1985) ‘Heads’. Journal of Linguistics 10. 1–30.